OpenStudy (anonymous):

It's been a while so can somebody show me how to figure this out? How do I find how many zeroes a function has over a set of complex numbers? (I'll post the specific function I need to work on in the comments)

OpenStudy (anonymous):

Okay well the equation thing isnt working for me so this is going to look really sloppy: f(x) = x^5 - 2x^4 + 8x^2 - 13x + 6

OpenStudy (anonymous):

I also need help on listing the possible rational zeroes, and factoring the polynomial completely over the set of complex numbers... but first things first.

OpenStudy (whpalmer4):

I worked this problem with someone else a few days ago: http://openstudy.com/study#/updates/52d71f53e4b0e63d49ac287c Unfortunately it will be a bit harder to read with OpenStudy not formatting the equations :-(

OpenStudy (anonymous):

Thank you! It's loading the equations for me just fine, it's just not letting me type them like I usually do.

OpenStudy (anonymous):

At the end, I need help completeing the square for x^2 - 2x + 3... I've never really been good at completeing squares

OpenStudy (whpalmer4):

Okay, to complete the square, we need to get the x^2 and x terms on one side and any solely numerical term on the other. x^2-2x = -3

OpenStudy (anonymous):

okay i understand that much c:

OpenStudy (whpalmer4):

Now we take half the value of the coefficient of the x term, square it, and add it to both sides.

OpenStudy (anonymous):

so, 1?

OpenStudy (whpalmer4):

Remember that (x+a)(x+a) = x^2 + ax + ax + a^2 = x^2 + 2ax + a^2 So we are just taking that middle term and making it equal to 2ax Dividing by 2 and squaring it gives us a^2

OpenStudy (anonymous):

So would it be x^2 - 2x + 1 and the factors be (x - 1)(x - 1)?

OpenStudy (anonymous):

and if we added 1 to the other side then that would be -2

OpenStudy (whpalmer4):

As you said, we get 1 after dividing by 2 and squaring. Now we add that to each side of the equation: x^2 -2x +1 = -3 + 1 We can rewrite the left side as a square: (x-1)(x-1) = 2 Take the square root of both sides and we get (X-1) = sqrt(2) But we need to use both the positive and negative square roots to get all the solutions

OpenStudy (whpalmer4):

Sorry, I'm being pestered for attention by a demanding little kid which s slowing down my typing :-)

OpenStudy (whpalmer4):

x-1 = sqrt(2) x = 1+sqrt(2) (Solution 1) x-1 = -sqrt(2) x = 1-sqrt(2) (Solution 2)

OpenStudy (anonymous):

That's fine, lol take your time. Oh so then two of the complex solutions are 1+sqrt(2) and 1-sqrt(2)?

OpenStudy (anonymous):

but wait that 2 was negative, right? we can't find the square root of negatives?

OpenStudy (whpalmer4):

Whoops, we dropped the minus sign on the 2! Those should be square root of -2, or we could keep it as I wrote it but multiply each radical by i

OpenStudy (whpalmer4):

1+i*sqrt(2) and 1-i*sqrt(2)

OpenStudy (anonymous):

ohhh okay I see

OpenStudy (whpalmer4):

I wish they would fix the %}%%## equation formatting

OpenStudy (whpalmer4):

Anyhow, do you understand completing the square now?

OpenStudy (anonymous):

yeah me too /: it's alright though, i still understand everything you said (: thank you very much! and yes, it makes a lot more sense to me now

OpenStudy (anonymous):

So the roots of this equation are (x - 1)(x - 1)(x + 2)(1 + i*sqrt(2))(1 - i*sqrt(2))?

OpenStudy (anonymous):

Is that how I would write it out factored?

OpenStudy (whpalmer4):

Close: (x-1)(x-1)(x+2)(x-1-i*sqrt(2))(x-1-i*sqrt(2)) assuming I didn't make any mistakes typing

OpenStudy (whpalmer4):

Uh, I did, the second one should have the other sign for the i*sqrt(2) part

OpenStudy (anonymous):

okay, should the last two be (x - 1 - i*sqrt(2)) or should one of them be (x - 1 + i*sqrt(2))?

OpenStudy (anonymous):

Oh okay (:

OpenStudy (anonymous):

Thank you very much!

OpenStudy (whpalmer4):

Yeah, that second one you wrote is correct